### v1

# ,  Wei Jin

Abstract
A subset of vertices is called a dissociation set if it induces a subgraph with vertex degree at most one. Recently, Yuan et al. established the upper bound of the maximum number of dissociation sets among all connected graphs of order $n$ and characterized the corresponding extremal graphs. They also proposed a question regarding the second-largest number of dissociation sets among all connected graphs of order $n$ and the corresponding extremal graphs. In this paper, we give a positive answer to this question.
Keywords: dissociation set; extremal enumeration; connected graph.
1 Introduction
We consider finite, simple, and undirected graphs and follow the terminology in [1]. The study of graph substructures and their enumeration has been a theme in extremal graph theory, dating back to a question posed by Erdős and Moser in 1960: what is the maximum possible number of maximal independent sets in general graphs of order $n$, and what are the graphs that achieve this maximum value? This was solved by Moon and Moser [6] in 1965. Since then, the maximum number of maximal or maximum independent sets on various graph families such as trees, unicyclic graphs, connected graphs and the graphs with at most $r$ cycles have been determined, refer [13, 3, 5, 7, 14, 18] for detail.
A subset $S$ of vertices in a graph $G$ is called a dissociation set if it induces a subgraph with vertex degree at most one. This concept, introduced by Yannakakis [15], not only generalizes the notions of independent sets and induced matchings but also forms a dual relationship with the 3-path vertex cover problem [2]. In recent years, scholars have begun to pay attention to determining the maximum (or minimum) number of maximal or maximum dissociation sets on various graph families and characterize the corroding extremal graphs such as trees [11, 10, 12], forests [8], unicyclic graphs [17], general and triangle-free graphs [9].
Recently, Yuan et al. [16] focused on the problem of the maximum number of all dissociation sets and the characterization of extremal graphs. They established the upper bound of the maximum number of dissociation sets among all connected graphs of order $n$ and characterized the corresponding extremal graphs. They also proposed a question regarding the second-largest number of dissociation sets and the corresponding extremal graphs as follows.
Question 1.1
Is $h(n)$ the second-largest number of dissociation sets among all connected graphs of order $n$ when $n\geq 10$? Do the extremal graphs belong to
$\begin{cases}K_{3}*\frac{n-3}{2}K_{2}\ or\ K_{1}*(2K_{1}\cup\frac{n-3}{2}K_{2})&\text{ if $n$ is odd; }\\ K_{3}*(\frac{n-4}{2}K_{2}\cup K_{1})\ or\ K_{1}*(3K_{1}\cup\frac{n-4}{2}K_{2})&\text{ if $n$ is even.}\end{cases}$
The operation * is a method of transforming several disjoint complete subgraphs into a connected graph, defined as follows. Let $G=K_{s_{1}}\cup K_{s_{2}}\cup\cdots\cup K_{s_{t}}$ be a disconnected graph whose all components are complete subgraphs, and $K_{r}$ a complete graph disjoint from $G$. $K_{r}*G$ is the graph obtained by selecting a vertex $u\in V(K_{r})$ and connecting $u$ to exactly one vertex from each component of $G$ (see Figure 1).
Figure 1: $K_{r}\ast(K_{s_{1}}\cup K_{s_{2}}\cup\cdots\cup K_{s_{t}})$.
In this paper, we give a positive answer to this question. The remainder of the paper is organized as follows: Section 2 introduces key definitions and preliminary lemmas; Section 3 presents the main results, including the upper bound for the second-largest number of dissociation sets and the corresponding extremal graphs.
2 Preliminaries
The set of all dissociation sets of $G$ is denoted by $\mathscr{D}(G)$, and the number of dissociation sets consisting of $k$ vertices is denoted by $d\left(G,k\right)$. For convenience, we define $d\left(G,0\right)=1$ for all graphs. The total number of dissociation sets $d(G)$ of $G$ is defined as
$d(G)=\sum_{k=0}^{\left|V(G)\right|}d(G,k).$
Let $G_{1},G_{2},\dots,G_{t}$ be all components of $G$. We can easily see that $d(G)=\prod_{i=1}^{i=t}d(G_{i})$. Two different vertices $u,v\in V(G)$ are called true twins if $N\left[u\right]=N\left[v\right]$ which indicates that $uv\in E(G)$ and are called false twins if $N(u)=N(v)$. It is easy to see that if $u,v$ are true twins in graph $G$, then $u,v$ are false twins in the subgraph $G-uv$.
A vertex with degree one is called a pendant vertex and the neighbor of a pendant vertex with degree at least two is called a quasi-pendant vertex. The dimension of cycle space is denoted by $c(G)=|E(G)|-|V(G)|+1$ where $|E(G)|$ and $|V(G)|$ represent the number of edges and vertices in $G$, respectively. Consider the cycle $C_{n}$ as $v_{1}v_{2}\cdots v_{n}v_{1}$. Denote by $C_{n}(k_{1},\cdots,k_{n})$ the unicyclic graph obtained from $C_{n}$ by identifying $v_{i}\in V(C_{n})$ with a vertex of pendant tree $T_{k_{i}}$ for $1\leq i\leq n$, where $k_{1}+\cdots+k_{n}$ equals to the order of $C_{n}(k_{1}\cdots k_{n})$.
Lemma 2.1
( [16]). Let $G$ be a graph of order $n$, then $d(G)\leq 2^{n}$, the equality holds if and only if $G\cong sK_{1}\cup tK_{2}$ where $s+2t=n$.
Lemma 2.2
( [16]). Let $G$ be a graph with $v\in V(G)$. Then
$d(G)=d(G-v)+d(G-N[v])+\sum_{u\in N(v)}d(G-N[u]\cup N[v])$
Lemma 2.3
( [16]). Let $G$ be a graph with $v\in V(G)$. Then $d(G)\leq d(G-uv)$, the equality holds if and only if $N_{G}[u]=N_{G}[v]$.
Theorem 2.4
( [16]). Let $G$ be a connected graph of order $n$. Then $d(G)\leq f(n)$, the equality holds if and only if $G\cong F_{n}$, where
$f(n):=\begin{cases}2^{n-1}+(n+3)\cdot 2^{\frac{n-5}{2}}&\text{if }n\text{ is odd};\\ 2^{n-1}+(n+6)\cdot 2^{\frac{n-6}{2}}&\text{if }n\text{ is even,}\end{cases}$
and
$F_{n}:=\begin{cases}K_{1}\ast\frac{n-1}{2}K_{2}&\text{ if $n$ is odd; }\\ P_{6}\ or\ K_{2}\ast 2K_{2}&\text{ if $n$ = 6; }\\ K_{2}\ast\frac{n-2}{2}K_{2}&\text{ if $n$ is even and $n$ $\neq$ 6.}\end{cases}$
Theorem 2.5
( [16]). Let $G$ be a unicyclic graph of order $n\geq 3$. Then $d(G)\leq h(n)$, the equality holds if and only if $G\cong U_{n}$, where
$h(n):=\begin{cases}2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}&\text{if }n\text{ is odd};\\ 2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}&\text{if }n\text{ is even and }n\neq 6;\\ 42&\text{if }n=6.\end{cases}$
and
$U_{n}:=\begin{cases}K_{3}\ast(\left\lfloor\frac{n-3}{2}\right\rfloor K_{2}\cup(n+1\bmod 2)K_{1})&\text{ if }n\neq 6;\\ K_{1}\ast(K_{3}\cup K_{2})&\text{ if }n=6.\end{cases}$
Lemma 2.6
Let $G$ be a graph with quasi-pendant vertex $u_{q}$. Suppose that $v_{1},v_{2},\cdots,v_{s}$ are pendant vertices that are adjacent to $u_{q}$ where $s\geq 2$. Then $d(G)<d(G+v_{s-1}v_{s}-u_{q}v_{s})$.
Proof. Note that $v_{s-1},v_{s}$ are true twins in $G+v_{s-1}v_{s}$. By Lemma 2.3, $d(G)=d(G+v_{s-1}v_{s})$. Note that $u_{q},v_{s}$ are not true twins in the graph $G+v_{s-1}v_{s}$. By Lemma 2.3, $d(G+v_{s-1}v_{s})<d(G+v_{s-1}v_{s}-u_{q}v_{s})$. Then $d(G)<d(G+v_{s-1}v_{s}-u_{q}v_{s})$. $\Box$
Remark. The proof of this lemma comes from [16], but its conclusion is stronger than the lemma above. For the sake of the completeness of the paper and to better serve this paper, we have rewritten the conclusion above.
Corollary 2.7
Let $T$ be a tree of order $n$ with the second-largest number of dissociation sets and $u_{q}$ a quasi-pendant vertex of $T$. If $u_{q}$ is connected to $s$ pendant vertices, then $s\leq 3$.
Proof. In contrast, suppose that $s\geq 4$. Choose four pendant vertices $v_{1},v_{2},v_{3},v_{4}$ that are adjacent to $u_{q}$. Note that $v_{1},v_{2}$ are true twins in $T+v_{1}v_{2}$. By Lemma 2.3, $d(T)=d(T+v_{1}v_{2})$. Let ${T}^{\prime}=T+v_{1}v_{2}-u_{q}v_{2}$ (see Figure 2). Note that $u_{q},v_{2}$ are not true twins in the graph $T+v_{1}v_{2}$. By Lemma 2.3, $d(T+v_{1}v_{2})<d(T+v_{1}v_{2}-u_{q}v_{2})=d(T^{\prime})$. Similarly, we can find that $d({T}^{\prime})=d({T}^{\prime}+v_{3}v_{4})<d({T}^{\prime}+v_{3}v_{4}-u_{q}v_{4})$. Let $T^{\prime\prime}={T}^{\prime}+v_{3}v_{4}-u_{q}v_{4}$. Consequently, $d(T)<d({T}^{\prime})<d(T^{\prime\prime})$ and $|V(T)|=|V(T^{\prime})|=|V(T^{\prime\prime})|$. This contradicts the fact that $T$ is the tree of order $n$ with the second-largest number of dissociation sets. So, $s\leq 3$. $\Box$
Figure 2: $T$, ${T}^{\prime}$, ${T}^{\prime\prime}$.
3 The second-largest number of dissociation sets in connected graph
In this section, we will give a positive answer to Question 1.1. In fact, we can replace the prerequisite $n\geq 10$ with $n\geq 9$. First, we determine the second-largest number of dissociation sets among all trees of order $n$ and characterize the structure of the corresponding extremal trees.
Lemma 3.1
[…]
Remark. The proof of this lemma is obtained by analyzing all trees of order $9$ or $10$. The conclusion is essential, but verification is unskilled and very cumbersome. So we provide the detailed proof in the appendix.
Theorem 3.2
Let $T$ be a tree of order $n(n\geq 9)$. If $T\ncong F_{n}$, then $d(T)\leq h(n)$, the equality holds if and only if $T\cong T_{n}$, where
$T_{n}:=\begin{cases}K_{1}\ast(2K_{1}\cup\frac{n-3}{2}K_{2}),&\text{ if $n$ is odd ;}\\ K_{1}\ast(3K_{1}\cup\frac{n-4}{2}K_{2}),&\text{ if $n$ is even ,}\end{cases}$
$F_{n}$ and $h(n)$ are defined in Theorem 2.4 and Theorem 2.5, respectively.
Proof. We shall prove by induction on $n$. By Lemma 3.1, the theorem holds for $n=9$ and $n=10$. Assume that the theorem is true when the order of the tree is less than $n$ where $n\geq 11$. Let $T$ be a tree of order $n$ such that $T\ncong F_{n}$. By Theorem 2.4, $d(T)<f(n)$. Let $T$ be a tree of order $n$ with the second-largest number of dissociation sets and choose a longest path $P=v_{0}v_{1}v_{2}\cdots v_{l}$ of $T$ starting from $v_{0}$. By Corollary 2.7, we can obtain ${\rm deg}_{T}(v_{1})\leq 4$. To be more specific, we have the following claim.
Claim 1. ${\rm deg}_{T}(v_{1})=2$.
Figure 3: Structure of $T$ in Claim 1.
On the contrary, suppose $3\leq{\rm deg}_{T}(v_{1})\leq 4$. By the choice of $P$, all components of $T-v_{1}$ are isolated vertices except the component that contains $v_{2}$. Let $u$ be a pendant vertex that is adjacent to $v_{1}$ where $u\not=v_{0}$. Let ${\widetilde{T}=T+v_{0}u-v_{1}u}$, see Figure 3. By Lemma 2.6, $d(T)<d(\widetilde{T})$. Furthermore, $\widetilde{T}\ncong F_{n}$ (by Theorem 2.4). It follows that $d(T)<d(\widetilde{T})<f(n)$. This is contradictory because $T$ is the tree with the second-largest number of dissociation sets. Thus, ${\rm deg}_{T}(v_{1})=2$.
Let $C$ be an arbitrary component of $T-v_{2}$ such that $V(P)\cap V(C)=\emptyset$. Then there are only two possible structures for $C$. Specifically, we have the following claim.
Claim 2: If $C$ is not an isolated vertex, then $C\cong K_{1,1}$.
In fact, suppose that $C$ is not an isolated vertex. By the choice of $P$, the diameter of the induced subgraph $T\left[V(C)\cup\left\{v_{2}\right\}\right]$ is no more than two. Thus, $C\cong K_{1,t}$. By Corollary 2.7, $t\leq 3$. To complete the proof of the claim, we need to show $t=1$. In fact, suppose $t\geq 2$. Choose a path $uvw$ from $C$. Clearly, $v$ is the center of $C$ which is adjacent to $v_{2}$ in $T$. By Lemma 2.6 and Theorem 2.4, $d(T)<d(T+wu-uv)<f(n)$. This contradicts the fact that $T$ is the tree with the second-largest number of dissociation sets. It follows that $C\cong K_{1,1}$.
Figure 4: Structure of $T$ in Claim 2.
Thus, every component of $T-v_{2}$ that does not contain $v_{3}$ is either an isolated vertex or an isolated edge. Suppose that there are $t_{1}$ isolated vertices and $t_{2}$ isolated edges in $T-v_{2}$, see Figure 4. By Corollary 2.7, $t_{1}\leq 3$. Let $T^{\prime}$ be the component of $T-v_{2}$ that contains $v_{3}$ and let $|V(T^{\prime})|=m$. Since $|V(T)|\geq 11$, $|V(T)|=t_{1}+2t_{2}+1+m\geq 11$. It follows that $2t_{2}+m\geq 7$. Thus, $m\geq 7$ or $t_{2}\geq 1$. Because $P=v_{0}v_{1}\cdots v_{l}$ is the longest path in $T$, $l\geq 4$. Specifically, we have the following claim.
Claim 3: $l=4$.
In contrast, suppose that $l>4$. Then $33\leq m\leq n-$ 3. By Lemma 2.2,
$d(T)$ $=d(T-v_{0})+d(T-v_{0}-v_{1})+d(T-v_{0}-v_{1}-v_{2}).$
By Theorem 2.4,
[…]
  • $d(T-v_{0})=f(n-1)$.
    By Theorem 2.4, $T-v_{0}\cong F_{n-1}$. It follows that $t_{1}=0$, $t_{2}=1$. Thus, there are only two possible structures for $T$ as shown in Figure 5. Suppose that $n$ is even ( $n\geq 11$). By Lemma 2.1,
    $d(T)$ $=d(T-v_{3})+d(T-N[v_{3}])+\sum_{u\in N(v_{3})}d(T-N[v_{3}]\cup N[u])$
    $=7\cdot 2^{n-4}+4\cdot 2^{\frac{n-4}{2}}+(\frac{n-4}{2}\cdot 4\cdot 2^{\frac{n-4}{2}-1}+2\cdot 2^{\frac{n-4}{2}})$
    $=7\cdot 2^{n-4}+(n+2)\cdot 2^{\frac{n-4}{2}}$
    $<h(n)=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}.$
    This is contradictory.
    Figure 5: The two possible structures of $T$ if $d(T-v_{0})=f(n-1)$.
    Suppose $n$ is odd ($n\geq 11$). By Lemma 2.1,
    $d(T)$ $=d(T-v_{3})+d(T-N[v_{3}])+\sum_{u\in N(v_{3})}d(T-N[v_{3}]\cup N[u])$
    $=2\cdot 7\cdot 2^{n-5}+4\cdot 2^{\frac{n-5}{2}}+(4\cdot 2^{\frac{n-5}{2}}+\frac{n-5}{2}\cdot 4\cdot 2^{\frac{n-5}{2}-1}+2\cdot 2^{\frac{n-5}{2}})$
    $=7\cdot 2^{n-4}+(n+5)\cdot 2^{\frac{n-5}{2}}$
    $<h(n)=2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}.$
    This is contradictory.
  • $d(T-v_{0})<f(n-1)$.
    Note that $T-v_{0}$ is connected and $d(T-v_{0})<f(n-1)$. By induction hypothesis, $d(T-v_{0})\leq h(n-1)$. Note that $T-v_{0}-v_{1}$ is connected. By Theorem 2.4, $d(T-v_{0}-v_{1})\leq f(n-2)$.
    Suppose that $n$ is even ($n\geq 11$). By Lemma 2.1,
    $d(T)$ $=d(T-v_{0})+d(T-v_{0}-v_{1})+d(T-v_{0}-v_{1}-v_{2})$
    $\leq h(n-1)+f(n-2)+7\cdot 2^{n-6}$
    $=[2^{n-2}+(n+8)\cdot 2^{\frac{n-8}{2}}]+[2^{n-3}+(n+4)\cdot 2^{\frac{n-8}{2}}]+7\cdot 2^{n-6}$
    $=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}-(2^{n-6}-n\cdot 2^{\frac{n-8}{2}})$
    $<h(n)=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}.$
    The last inequality follows from the fact that $02^{n-6}-n\cdot 2^{\frac{n-8}{2}}>$ 0 when $n\geq 12$. This is contradictory.
    Suppose that $n$ is odd ($n\geq 11$). By Lemma 2.1,
    $d(T)$ $=d(T-v_{0})+d(T-v_{0}-v_{1})+d(T-v_{0}-v_{1}-v_{2})$
    $\leq h(n-1)+f(n-2)+7\cdot 2^{n-6}$
    $=[2^{n-2}+(n+11)\cdot 2^{\frac{n-9}{2}}]+[2^{n-3}+(n+1)\cdot 2^{\frac{n-7}{2}}]+7\cdot 2^{n-6}$
    $=2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}-(2^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}})$
    $<h(n)=2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}.$
    The last inequality follows from the fact that $02^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}}>$ 0 when $n\geq 11$. This is contradictory. Here, we have completed the proof of Claim $3$.
By Claim $3$, the component of $T-v_{2}$ that contains $v_{3}$ is a star $K_{1,x}$. Relabel the vertices of the path $P=v_{0}v_{1}\cdots v_{l}$ as $u_{0}u_{1}\cdots u_{l}$ where $u_{i}=v_{l-i}$ for all $i=0,1,\cdots,l$. By Claim $1$, $x=1$. Thus, every component of $T-v_{2}$ is $K_{1}$ or $K_{2}$. It follows that
$T\in\{K_{1}\ast\frac{n-1}{2}K_{2},K_{2}\ast\frac{n-2}{2}K_{2},K_{2}\ast(\frac{n-3}{2}K_{2}\cup K_{1}),K_{2}\ast(\frac{n-4}{2}K_{2}\cup 2K_{1})\}$(see Figure 6),
the integer $n$ in the first and third elements of the set above are odd, and the other two are even. By Theorem 2.4, $K_{1}\ast\frac{n-1}{2}K_{2}$ and $K_{2}\ast\frac{n-2}{2}K_{2}$ are two trees with the largest number of dissociation sets. Hence, $T\in\{K_{2}\ast(\frac{n-3}{2}K_{2}\cup K_{1}),K_{2}\ast(\frac{n-4}{2}K_{2}\cup 2K_{1})\}$, The integer $n$ in the first element of the set is odd, and the second is even. $\Box$
Figure 6: The possible structure of $T$.
Theorem 3.3
Let $G$ be a connected graph of order $n\geq 9$. If $G\ncong F_{n}$, then $d(G)\leq h(n)$, the equality holds if and only if $G\in\left\{U_{n},T_{n}\right\}$, where $F_{n}$, $U_{n}$, $h(n)$, and $T_{n}$ are defined in Theorem 2.4, Theorem 2.5, and Theorem 3.2 respectively.
Proof. Let $c(G)$ be the dimension of cycle spaces of $G$. If $c(G)\leq 1$, then $G$ is a tree or a unicyclic graph. By Theorem 2.5 and Theorem 3.2, the result holds. To complete the proof, we just need to show that $d(G)<h(n)$ if $c(G)=d\geq 2$.
In fact, suppose $d\geq 2$. Since $G$ is connected, we can choose $d-1$ edges $u_{1}v_{1},u_{2}v_{2},\ldots,u_{d-1}v_{d-1}$ of $G$ such that the subgraph
$G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-1}v_{d-1}:={G}^{\prime}$
is a unicyclic graph. Then, we get
$d(G)$ $\leq d(G-u_{1}v_{1})$
[…]
$\leq d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})$
$\leq d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2}-u_{d-1}v_{d-1})$
$=d({G}^{\prime})\leq h(n).$
Hence, $d(G)=h(n)$ if and only if all of the above inequalities hold. By Theorem 2.5, $d({G}^{\prime})=h(n)$ is equivalent to ${G}^{\prime}\cong U_{n}$. By Lemma 2.3, $d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})=d({G}^{\prime})$ is equivalent to $u_{d-1},v_{d-1}$ are true twins in $G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2}$ It is follows that $u_{d-1},v_{d-1}$ are false twins in ${G}^{\prime}$. But this contradicts the fact that there are no false twins in ${G}^{\prime}\cong U_{n}$, where $U_{n}$ is defined in Theorem 2.5. Then $d(G)<h(n)$. $\Box$
Appendix
In the appendix, we will prove Lemma 3.1. When $n=9$, there are $47$ different trees in the sense of isomorphism, we label these trees as $T_{1}^{9},T_{2}^{9},\cdots T_{47}^{9}$ in order, This figure is from page 233 of [4]. see Figure 7.
Figure 7: When $n=9$, 47 structures of tree $T$.
By Lemma 2.6, it is easy to see that
$\begin{array}[]{llll}d(T_{2}^{9})<d(T_{1}^{9}),&d(T_{7}^{9})<d(T_{3}^{9}),&d(T_{8}^{9})<d(T_{5}^{9}),&d(T_{9}^{9})<d(T_{6}^{9}),\vspace{1.0ex}\\ d(T_{13}^{9})<d(T_{10}^{9}),&d(T_{19}^{9})<d(T_{1}^{9}),&d(T_{20}^{9})<d(T_{3}^{9}),&d(T_{21}^{9})<d(T_{5}^{9}),\vspace{1.0ex}\\ d(T_{23}^{9})<d(T_{4}^{9}),&d(T_{24}^{9})<d(T_{4}^{9}),&d(T_{25}^{9})<d(T_{6}^{9}),&d(T_{28}^{9})<d(T_{3}^{9}),\vspace{1.0ex}\\ d(T_{29}^{9})<d(T_{5}^{9}),&d(T_{30}^{9})<d(T_{22}^{9}),&d(T_{31}^{9})<d(T_{6}^{9}),&d(T_{32}^{9})<d(T_{10}^{9}),\vspace{1.0ex}\\ d(T_{33}^{9})<d(T_{22}^{9}),&d(T_{36}^{9})<d(T_{10}^{9}),&d(T_{38}^{9})<d(T_{27}^{9}),&d(T_{39}^{9})<d(T_{27}^{9}),\vspace{1.0ex}\\ d(T_{40}^{9})<d(T_{26}^{9}),&d(T_{41}^{9})<d(T_{27}^{9}),&d(T_{43}^{9})<d(T_{26}^{9}),&d(T_{44}^{9})<d(T_{4}^{9}),\vspace{1.0ex}\\ d(T_{45}^{9})<d(T_{6}^{9}),&d(T_{46}^{9})<d(T_{26}^{9}),&d(T_{47}^{9})<d(T_{6}^{9}).\end{array}$
By Corollary 2.7, $T_{12}^{9}$, $T_{15}^{9}$, $T_{16}^{9}$, $T_{17}^{9}$, $T_{18}^{9}$, $T_{34}^{9}$, $T_{35}^{9}$, $T_{37}^{9}$, $T_{42}^{9}$ are not the trees with the second-largest number of dissociation sets. By Lemma 2.2 , Theorem 2.4 and Theorem 2.5, we have $d(T_{1}^{9})=274<h(9)=292$, $d(T_{11}^{9})=f(9)=304$, $d(T_{14}^{9})=h(9)=292$. Therefore, the trees with the second-largest number of dissociation sets belong to $\{T_{3}^{9},T_{4}^{9},T_{5}^{9},T_{6}^{9},T_{10}^{9},T_{22}^{9},T_{26}^{9},T_{27}^{9}\}$.
By Lemma 2.2, $d(P_{n})=d(P_{n-1})+d(P_{n-2})+d(P_{n-3})$. It can be easily calculated that $d(P_{1})=2$, $d(P_{2})=4$, $d(P_{3})=7$, $d(P_{4})=13$, $d(P_{5})=24$, $d(P_{6})=44$, $d(P_{7})=81$, $d(P_{8})=149$, $d(P_{9})=274$. By Lemma 2.2 , we can calculate the number of dissociation sets of these eight graphs as follows by choosing $v_{0}$ as shown in Figure 8.
In fact,
$d(T_{3}^{9})$ $=d(T_{3}^{9}-v_{0})+d(T_{3}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{3}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{2})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{1})\cdot d(P_{4})+d(P_{4})+d(P_{1})\cdot d(P_{3})$
$=271<h(9);$
$d(T_{4}^{9})$ $=d(T_{4}^{9}-v_{0})+d(T_{4}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{4}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{3})\cdot d(P_{4})\cdot d(P_{1})+2\cdot d(P_{2})\cdot d(P_{3})+d(P_{1})\cdot d(P_{3})+d(P_{2})^{2}$
$=268<h(9);$
$d(T_{5}^{9})$ $=d(T_{5}^{9}-v_{0})+d(T_{5}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{5}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{2})^{2}\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})+d(P_{1})\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})$
$=280<h(9);$
$d(T_{6}^{9})$ $=d(T_{6}^{9}-v_{0})+d(T_{6}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{6}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{2})\cdot d(P_{3})^{2}+d(P_{1})\cdot d(P_{2})^{2}+d(P_{2})\cdot d(P_{1})^{2}\cdot 2+d(P_{2})^{2}$
$=276<h(9);$
$d(T_{10}^{9})$ $=d(T_{10}^{9}-v_{0})+d(T_{10}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{10}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{2})^{2}\cdot d(P_{3})+d(P_{2})\cdot d(P_{1})^{2}\cdot 2+2\cdot d(P_{1})\cdot d(P_{2})+d(P_{1})^{3}$
$=280<h(9);$
$d(T_{22}^{9})$ $=d(T_{22}^{9}-v_{0})+d(T_{22}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{22}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{4})^{2}+d(P_{1})^{2}\cdot d(P_{2})^{2}+2\cdot d(P_{1})^{2}\cdot d(P_{2})$
$=265<h(9);$
$d(T_{26}^{9})$ $=d(T_{26}^{9}-v_{0})+d(T_{26}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{26}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{2})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{1})^{2}\cdot d(P_{3})+d(P_{1})\cdot d(P_{3})+d(P_{1})\cdot d(P_{2})$
$=270<h(9);$
$d(T_{27}^{9})$ $=d(T_{27}^{9}-v_{0})+d(T_{27}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{27}^{9}-N[v_{0}]\cup N[u])$
$=d(P_{2})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{1})\cdot d(P_{2})^{2}+d(P_{2})^{2}+d(P_{1})^{3}$
$=280<h(9).$
Figure 8: Eight trees that may have the second-largest number of dissociation sets.
So, when $n=9$, the conclusion is true.
When $n=10$, there are $106$ different trees in the sense of isomorphism, we label these trees as $T_{1}^{10},T_{2}^{10}$, $\cdots$, $T_{106}^{10}$ in order, §§§This figure is from page 234 of [4]. see Figure 9. By Lemma 2.6, it is easy to see that
$\begin{array}[]{llll}d(T_{2}^{10})<d(T_{1}^{10}),&d(T_{9}^{10})<d(T_{3}^{10}),&d(T_{10}^{10})<d(T_{5}^{10}),&d(T_{11}^{10})<d(T_{7}^{10}),\vspace{1.0ex}\\ d(T_{16}^{10})<d(T_{12}^{10}),&d(T_{17}^{10})<d(T_{13}^{10}),&d(T_{18}^{10})<d(T_{14}^{10}),&d(T_{27}^{10})<d(T_{1}^{10}),\vspace{1.0ex}\\ d(T_{28}^{10})<d(T_{3}^{10}),&d(T_{29}^{10})<d(T_{5}^{10}),&d(T_{31}^{10})<d(T_{4}^{10}),&d(T_{32}^{10})<d(T_{6}^{10}),\vspace{1.0ex}\\ d(T_{33}^{10})<d(T_{7}^{10}),&d(T_{36}^{10})<d(T_{4}^{10}),&d(T_{37}^{10})<d(T_{7}^{10}),&d(T_{39}^{10})<d(T_{8}^{10}),\vspace{1.0ex}\\ d(T_{44}^{10})<d(T_{3}^{10}),&d(T_{45}^{10})<d(T_{30}^{10}),&d(T_{46}^{10})<d(T_{5}^{10}),&d(T_{47}^{10})<d(T_{34}^{10}),\vspace{1.0ex}\\ d(T_{48}^{10})<d(T_{7}^{10}),&d(T_{49}^{10})<d(T_{35}^{10}),&d(T_{50}^{10})<d(T_{35}^{10}),&d(T_{51}^{10})<d(T_{12}^{10}),\vspace{1.0ex}\\ d(T_{52}^{10})<d(T_{38}^{10}),&d(T_{53}^{10})<d(T_{7}^{10}),&d(T_{54}^{10})<d(T_{41}^{10}),&d(T_{55}^{10})<d(T_{42}^{10}),\vspace{1.0ex}\\ d(T_{56}^{10})<d(T_{13}^{10}),&d(T_{57}^{10})<d(T_{43}^{10}),&d(T_{58}^{10})<d(T_{41}^{10}),&d(T_{60}^{10})<d(T_{14}^{10}),\vspace{1.0ex}\\ d(T_{63}^{10})<d(T_{12}^{10}),&d(T_{65}^{10})<d(T_{13}^{10}),&d(T_{67}^{10})<d(T_{59}^{10}),&d(T_{68}^{10})<d(T_{14}^{10}),\vspace{1.0ex}\\ d(T_{73}^{10})<d(T_{30}^{10}),&d(T_{74}^{10})<d(T_{35}^{10}),&d(T_{75}^{10})<d(T_{59}^{10}),&d(T_{76}^{10})<d(T_{43}^{10}),\vspace{1.0ex}\\ d(T_{77}^{10})<d(T_{41}^{10}),&d(T_{80}^{10})<d(T_{59}^{10}),&d(T_{83}^{10})<d(T_{8}^{10}),&d(T_{84}^{10})<d(T_{41}^{10}),\vspace{1.0ex}\\ d(T_{85}^{10})<d(T_{40}^{10}),&d(T_{86}^{10})<d(T_{42}^{10}),&d(T_{88}^{10})<d(T_{7}^{10}),&d(T_{89}^{10})<d(T_{34}^{10}),\vspace{1.0ex}\\ d(T_{90}^{10})<d(T_{38}^{10}),&d(T_{91}^{10})<d(T_{4}^{10}),&d(T_{92}^{10})<d(T_{6}^{10}),&d(T_{93}^{10})<d(T_{13}^{10}),\vspace{1.0ex}\\ d(T_{94}^{10})<d(T_{41}^{10}),&d(T_{95}^{10})<d(T_{7}^{10}),&d(T_{96}^{10})<d(T_{41}^{10}),&d(T_{97}^{10})<d(T_{42}^{10}),\vspace{1.0ex}\\ d(T_{98}^{10})<d(T_{87}^{10}),&d(T_{99}^{10})<d(T_{34}^{10}),&d(T_{100}^{10})<d(T_{38}^{10}),&d(T_{101}^{10})<d(T_{13}^{10}),\vspace{1.0ex}\\ d(T_{102}^{10})<d(T_{87}^{10}),&d(T_{103}^{10})<d(T_{41}^{10}),&d(T_{104}^{10})<d(T_{8}^{10}),&d(T_{105}^{10})<d(T_{40}^{10}).\end{array}$
Figure 9: When $n=10$, 106 structures of tree $T$.
By Corollary 2.7, $T_{15}^{10}$, $T_{20}^{10}$, $T_{21}^{10}$, $T_{23}^{10}$, $T_{24}^{10}$, $T_{25}^{10}$, $T_{26}^{10}$, $T_{61}^{10}$, $T_{62}^{10}$, $T_{64}^{10}$, $T_{66}^{10}$, $T_{69}^{10}$, $T_{70}^{10}$, $T_{71}^{10}$, $T_{72}^{10}$, $T_{78}^{10}$, $T_{79}^{10}$, $T_{81}^{10}$, $T_{82}^{10}$, $T_{106}^{10}$ are not the trees with the second-largest number of dissociation sets. By Lemma 2.2 , Theorem 2.4 and Theorem 2.5, we have $d(T_{1}^{10})=504<h(10)=556$, $d(T_{19}^{10})=f(10)=576$, $d(T_{22}^{10})=h(10)=556$. Therefore, the graphs with the second-largest number of dissociation sets belong to
$\{T_{3}^{10}$, $T_{4}^{10}$, $T_{5}^{10}$, $T_{6}^{10}$, $T_{7}^{10}$, $T_{8}^{10}$, $T_{12}^{10}$, $T_{13}^{10}$, $T_{14}^{10}$, $T_{30}^{10}$, $T_{34}^{10}$, $T_{35}^{10}$, $T_{38}^{10}$, $T_{40}^{10}$, $T_{41}^{10}$, $T_{42}^{10}$, $T_{43}^{10}$, $T_{59}^{10}$, $T_{87}^{10}\}$.
By Lemma 2.2 , we can calculate the number of dissociation sets of these nineteen trees by choosing $v_{0}$ as shown in Figure 10.
$d(T_{3}^{10})$ $=d(T_{3}^{10}-v_{0})+d(T_{3}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{3}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{2})\cdot d(P_{1})\cdot d(P_{6})+2\cdot d(P_{1})\cdot d(P_{5})+d(P_{5})+d(P_{1})\cdot d(P_{4})$
$=498<h(10);$
$d(T_{4}^{10})$ $=d(T_{4}^{10}-v_{0})+d(T_{4}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{4}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{3})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{2})\cdot d(P_{4})+d(P_{1})\cdot d(P_{4})+d(P_{2})\cdot d(P_{3})$
$=494<h(10);$
$d(T_{5}^{10})$ $=d(T_{5}^{10}-v_{0})+d(T_{5}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{5}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{2})^{2}\cdot d(P_{5})+d(P_{1})^{2}\cdot d(P_{4})+2\cdot d(P_{1})\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})$
$=516<h(10);$
$d(T_{6}^{10})$ $=d(T_{6}^{10}-v_{0})+d(T_{6}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{6}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{4})^{2}\cdot d(P_{1})+d(P_{3})^{2}\cdot 2+2\cdot d(P_{2})\cdot d(P_{3})$
$=492<h(10);$
$d(T_{7}^{10})$ $=d(T_{7}^{10}-v_{0})+d(T_{7}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{7}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{2})\cdot d(P_{7})+d(P_{1})\cdot d(P_{2})\cdot d(P_{4})+d(P_{2})\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})$
$=508<h(10);$
$d(T_{8}^{10})$ $=d(T_{8}^{10}-v_{0})+d(T_{8}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{8}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{3})^{3}+d(P_{2})^{3}+3\cdot d(P_{1})\cdot d(P_{2})^{2}$
$=503<h(10);$
$d(T_{12}^{10})$ $=d(T_{12}^{10}-v_{0})+d(T_{12}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{12}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{2})^{2}\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})\cdot 2+d(P_{1})\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})$
$=516<h(10);$
$d(T_{13}^{10})$ $=d(T_{13}^{10}-v_{0})+d(T_{13}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{13}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{3})^{2}\cdot d(P_{2})+d(P_{2})^{2}\cdot d(P_{1})\cdot 2+d(P_{2})^{2}+d(P_{1})^{2}\cdot d(P_{2})\cdot 2$
$=504<h(10);$
$d(T_{14}^{10})$ $=d(T_{14}^{10}-v_{0})+d(T_{14}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{14}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{2})^{3}\cdot d(P_{3})+d(P_{1})^{3}\cdot d(P_{2})+d(P_{1})^{2}\cdot d(P_{2})\cdot 3+d(P_{1})^{4}$
$=544<h(10);$
Figure 10: Nineteen trees that may have the second-largest number of dissociation sets.
$d(T_{30}^{10})$ $=d(T_{30}^{10}-v_{0})+d(T_{30}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{30}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{4})\cdot 23+d(P_{1})\cdot d(P_{2})\cdot d(P_{4})+d(P_{1})\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{2})^{2}$
$=493<h(10);$
$d(T_{34}^{10})$ $=d(T_{34}^{10}-v_{0})+d(T_{34}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{34}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{4})\cdot d(P_{5})+d(P_{1})^{2}\cdot d(P_{2})\cdot d(P_{3})+d(P_{1})^{2}\cdot d(P_{3})+d(P_{1})\cdot d(P_{2})^{2}$
$=484<h(10);$
$d(T_{35}^{10})$ $=d(T_{35}^{10}-v_{0})+d(T_{35}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{35}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{4})\cdot d(P_{5})+d(P_{1})\cdot d(P_{2})^{3}+d(P_{1})\cdot d(P_{2})^{2}+d(P_{1})^{3}\cdot d(P_{2})$
$=504<h(10);$
$d(T_{38}^{10})$ $=d(T_{38}^{10}-v_{0})+d(T_{38}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{38}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{4})^{2}+d(P_{1})\cdot d(P_{2})\cdot d(P_{3})\cdot 2+d(P_{1})\cdot d(P_{3})+d(P_{2})^{2}\cdot d(P_{1})$
$=496<h(10);$
$d(T_{40}^{10})$ $=d(T_{40}^{10}-v_{0})+d(T_{40}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{40}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{3})\cdot d(P_{5})+d(P_{1})\cdot d(P_{2})\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{3})+d(P_{2})^{2}$
$=492<h(10);$
$d(T_{41}^{10})$ $=d(T_{41}^{10}-v_{0})+d(T_{41}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{41}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{2})\cdot d(P_{6})+d(P_{1})\cdot d(P_{2})\cdot d(P_{3})\cdot 2+d(P_{2})\cdot d(P_{3})+d(P_{1})^{2}\cdot d(P_{2})$
$=508<h(10);$
$d(T_{42}^{10})$ $=d(T_{42}^{10}-v_{0})+d(T_{42}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{42}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{2})^{2}\cdot d(P_{5})+d(P_{1})^{3}\cdot d(P_{3})+d(P_{1})^{2}\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})$
$=512<h(10);$
$d(T_{43}^{10})$ $=d(T_{43}^{10}-v_{0})+d(T_{43}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{43}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{2})^{2}\cdot d(P_{5})+d(P_{1})^{2}\cdot d(P_{2})^{2}+d(P_{1})\cdot d(P_{2})^{2}\cdot 2+d(P_{1})^{4}$
$=528<h(10);$
$d(T_{59}^{10})$ $=d(T_{59}^{10}-v_{0})+d(T_{59}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{59}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{2})^{2}\cdot d(P_{4})+d(P_{1})^{3}\cdot d(P_{2})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})\cdot 2+d(P_{1})^{3}$
$=520<h(10);$
$d(T_{87}^{10})$ $=d(T_{87}^{10}-v_{0})+d(T_{87}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{87}^{10}-N[v_{0}]\cup N[u])$
$=d(P_{1})\cdot d(P_{4})^{2}+d(P_{1})^{2}\cdot d(P_{2})^{2}\cdot 2+d(P_{1})^{2}\cdot d(P_{2})\cdot 2$
$=498<h(10).$
[…]

### v2

@@
 This was solved by Moon and Moser [6] in 1965.
-Since then, the maximum number of maximal or maximum independent sets on various graph families such as trees, unicyclic graphs, connected graphs and the graphs with at most $r$ cycles have been determined, refer [13, 3, 5, 7, 14, 18] for detail.
+Since then, the maximum number of maximal or maximum independent sets on various graph families such as trees, unicyclic graphs, connected graphs and the graphs with at most $r$ cycles have been determined ; see [13, 3, 5, 7, 14, 18] for details.
 A subset $S$ of vertices in a graph $G$ is called a dissociation set if it induces a subgraph with vertex degree at most one.
 This concept, introduced by Yannakakis [15], not only generalizes the notions of independent sets and induced matchings but also forms a dual relationship with the 3-path vertex cover problem [2].
-In recent years, scholars have begun to pay attention to determining the maximum (or minimum) number of maximal or maximum dissociation sets on various graph families and characterize the corroding extremal graphs such as trees [11, 10, 12], forests [8], unicyclic graphs [17], general and triangle-free graphs [9].
+In recent years, scholars have begun to pay attention to determining the maximum (or minimum) number of maximal or maximum dissociation sets on various graph families and characterize the corresponding extremal graphs such as trees [11, 10, 12], forests [8], unicyclic graphs [17], general and triangle-free graphs [9].
 Recently, Yuan et al.
 [16] focused on the problem of the maximum number of all dissociation sets and the characterization of extremal graphs.
-They established the upper bound of the maximum number of dissociation sets among all connected graphs of order $n$ and characterized the corresponding extremal graphs.
-They also proposed a question regarding the second-largest number of dissociation sets and the corresponding extremal graphs as follows.
+They established the maximum number of dissociation sets among all connected graphs of order $n$ and characterized the corresponding extremal graphs.
+They also proposed a question of determining the second-largest number of dissociation sets and the corresponding extremal graphs as follows.
 Question 1.1
 Is $h(n)$ the second-largest number of dissociation sets among all connected graphs of order $n$ when $n\geq 10$? Do the extremal graphs belong to
-$\begin{cases}K_{3}*\frac{n-3}{2}K_{2}\ or\ K_{1}*(2K_{1}\cup\frac{n-3}{2}K_{2})&\text{ if $n$ is odd; }\\ K_{3}*(\frac{n-4}{2}K_{2}\cup K_{1})\ or\ K_{1}*(3K_{1}\cup\frac{n-4}{2}K_{2})&\text{ if $n$ is even.}\end{cases}$
+[…]
+Here, $h(n)$ is defined as follows
+[…]
 The operation * is a method of transforming several disjoint complete subgraphs into a connected graph, defined as follows.
@@
 In this paper, we give a positive answer to this question.
-The remainder of the paper is organized as follows: Section 2 introduces key definitions and preliminary lemmas; Section 3 presents the main results, including the upper bound for the second-largest number of dissociation sets and the corresponding extremal graphs.
+The remainder of the paper is organized as follows: Section 2 introduces key definitions and preliminary lemmas; Section 3 presents the main results, including the second-largest number of dissociation sets in connected graphs of order $n$ and the corresponding extremal graphs.
 2 Preliminaries
@@
 Two different vertices $u,v\in V(G)$ are called true twins if $N\left[u\right]=N\left[v\right]$ which indicates that $uv\in E(G)$ and are called false twins if $N(u)=N(v)$.
-It is easy to see that if $u,v$ are true twins in graph $G$, then $u,v$ are false twins in the subgraph $G-uv$.
+It is easy to see that if $u$ and $v$ are true twins in graph $G$, then $u$ and $v$ are false twins in the subgraph GuvG-uv.
 A vertex with degree one is called a pendant vertex and the neighbor of a pendant vertex with degree at least two is called a quasi-pendant vertex.
-The dimension of cycle space is denoted by $c(G)=|E(G)|-|V(G)|+1$ where $|E(G)|$ and $|V(G)|$ represent the number of edges and vertices in $G$, respectively.
+The dimension of cycle space of a connected graph $G$ is denoted by c(G)=|E(G)||V(G)|+1c(G)=|E(G)|-|V(G)|+1 where $|E(G)|$ and $|V(G)|$ represent the number of edges and vertices in $G$, respectively.
 Consider the cycle $C_{n}$ as $v_{1}v_{2}\cdots v_{n}v_{1}$.
-Denote by $C_{n}(k_{1},\cdots,k_{n})$ the unicyclic graph obtained from $C_{n}$ by identifying $v_{i}\in V(C_{n})$ with a vertex of pendant tree $T_{k_{i}}$ for $1\leq i\leq n$, where $k_{1}+\cdots+k_{n}$ equals to the order of $C_{n}(k_{1}\cdots k_{n})$.
+Denote by $C_{n}(k_{1},\cdots,k_{n})$ the unicyclic graph obtained from $C_{n}$ by identifying $v_{i}\in V(C_{n})$ with a vertex of pendant tree $T_{k_{i}}$ for $1\leq i\leq n$, where $k_{1}+\cdots+k_{n}$ equals to the order of $C_{n}(k_{1},\cdots,k_{n})$.
 Lemma 2.1
@@
 ( [16]). Let $G$ be a graph with $v\in V(G)$. Then
-$d(G)=d(G-v)+d(G-N[v])+\sum_{u\in N(v)}d(G-N[u]\cup N[v])$
+$d(G)=d(G-v)+d(G-N[v])+\sum_{u\in N(v)}d(G-N[u]\cup N[v]){\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}.}$
 Lemma 2.3
-( [16]). Let $G$ be a graph with $v\in V(G)$. Then $d(G)\leq d(G-uv)$, the equality holds if and only if $N_{G}[u]=N_{G}[v]$.
+( [16]). Let $G$ be a graph with $uv\in E(G)$. Then $d(G)\leq d(G-uv)$, the equality holds if and only if $N_{G}[u]=N_{G}[v]$.
 Theorem 2.4
@@
 and
-$F_{n}:=\begin{cases}K_{1}\ast\frac{n-1}{2}K_{2}&\text{ if $n$ is odd; }\\ P_{6}\ or\ K_{2}\ast 2K_{2}&\text{ if $n$ = 6; }\\ K_{2}\ast\frac{n-2}{2}K_{2}&\text{ if $n$ is even and $n$ $\neq$ 6.}\end{cases}$
+[…]
 Theorem 2.5
 ( [16]). Let $G$ be a unicyclic graph of order $n\geq 3$. Then $d(G)\leq h(n)$, the equality holds if and only if $G\cong U_{n}$, where
-$h(n):=\begin{cases}2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}&\text{if }n\text{ is odd};\\ 2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}&\text{if }n\text{ is even and }n\neq 6;\\ 42&\text{if }n=6.\end{cases}$
+[…]
 and
@@
 Lemma 2.6
-Let $G$ be a graph with quasi-pendant vertex $u_{q}$.
+Let $G$ be a graph of order $n\geq 4$ with quasi-pendant vertex $u_{q}$.
 Suppose that $v_{1},v_{2},\cdots,v_{s}$ are pendant vertices that are adjacent to $u_{q}$ where $s\geq 2$.
@@
 Proof.
-Note that $v_{s-1},v_{s}$ are true twins in $G+v_{s-1}v_{s}$.
+Note that $v_{s-1}$ and $v_{s}$ are true twins in $G+v_{s-1}v_{s}$.
 By Lemma 2.3, $d(G)=d(G+v_{s-1}v_{s})$.
-Note that $u_{q},v_{s}$ are not true twins in the graph $G+v_{s-1}v_{s}$.
+Note that $u_{q}$ and $v_{s}$ are not true twins in the graph $G+v_{s-1}v_{s}$.
 By Lemma 2.3, $d(G+v_{s-1}v_{s})<d(G+v_{s-1}v_{s}-u_{q}v_{s})$.
@@
 Choose four pendant vertices $v_{1},v_{2},v_{3},v_{4}$ that are adjacent to $u_{q}$.
-Note that $v_{1},v_{2}$ are true twins in $T+v_{1}v_{2}$.
-By Lemma 2.3, $d(T)=d(T+v_{1}v_{2})$.
-Let ${T}^{\prime}=T+v_{1}v_{2}-u_{q}v_{2}$ (see Figure 2).
-Note that $u_{q},v_{2}$ are not true twins in the graph $T+v_{1}v_{2}$.
-By Lemma 2.3, $d(T+v_{1}v_{2})<d(T+v_{1}v_{2}-u_{q}v_{2})=d(T^{\prime})$.
-Similarly, we can find that $d({T}^{\prime})=d({T}^{\prime}+v_{3}v_{4})<d({T}^{\prime}+v_{3}v_{4}-u_{q}v_{4})$.
-Let $T^{\prime\prime}={T}^{\prime}+v_{3}v_{4}-u_{q}v_{4}$.
+By Lemma 2.6, $d(T)<d(T+v_{1}v_{2}-u_{q}v_{2}):=d(T^{\prime})$.
+Similarly, $d({T}^{\prime})<d({T}^{\prime}+v_{3}v_{4}-u_{q}v_{4}):=d(T^{\prime\prime})$.
 Consequently, $d(T)<d({T}^{\prime})<d(T^{\prime\prime})$ and $|V(T)|=|V(T^{\prime})|=|V(T^{\prime\prime})|$.
@@
 $\Box$
-Figure 2: $T$, ${T}^{\prime}$, ${T}^{\prime\prime}$.
+Figure 2: $T$, $^{\prime}, T′′{T}^{\prime\prime}$.
 3 The second-largest number of dissociation sets in connected graph
@@
 Lemma 3.1
-[…]
+Let $T$ be a tree of order n(n=9or 10)n(n=9\ or\ 10) where $T\notin\{K_{1}\ast 4K_{2},K_{2}\ast 4K_{2}\}$.
+Then $d(T)\leq h(|V(T)|)$, the equality holds if and only if $T\in\{K_{1}\ast(2K_{1}\cup 3K_{2}),K_{1}\ast(3K_{1}\cup 3K_{2})\}$ where $h(9)=292,h(10)=556$.
 Remark.
@@
 Let $T$ be a tree of order $n(n\geq 9)$. If $T\ncong F_{n}$, then $d(T)\leq h(n)$, the equality holds if and only if $T\cong T_{n}$, where
-$T_{n}:=\begin{cases}K_{1}\ast(2K_{1}\cup\frac{n-3}{2}K_{2}),&\text{ if $n$ is odd ;}\\ K_{1}\ast(3K_{1}\cup\frac{n-4}{2}K_{2}),&\text{ if $n$ is even ,}\end{cases}$
+[…]
 $F_{n}$ and $h(n)$ are defined in Theorem 2.4 and Theorem 2.5, respectively.
@@
 Let $\widetilde{T}=T+v_{0}u-v_{1}u$}, see Figure 3.
-By Lemma 2.6, $d(T)<d(\widetilde{T})$.
+By Lemma 2.6, d(T)<d(T~)d(T)<d($\widetilde${T}).
 Furthermore, $\widetilde{T}\ncong F_{n}$ (by Theorem 2.4).
@@
 In fact, suppose $t\geq 2$.
-Choose a path $uvw$ from $C$.
+Choose a path uvwuvw from $C$.
 Clearly, $v$ is the center of $C$ which is adjacent to $v_{2}$ in $T$.
-By Lemma 2.6 and Theorem 2.4, $d(T)<d(T+wu-uv)<f(n)$.
+By Lemma 2.6 and Theorem 2.4, d(T)<d(T+wuuv)<f(n)d(T)<d(T+wu-uv)<f(n).
 This contradicts the fact that $T$ is the tree with the second-largest number of dissociation sets.
@@
 Claim 3: $l=4$.
-In contrast, suppose that $l>4$. Then $33\leq m\leq n-$ 3. By Lemma 2.2,
+In contrast, suppose that $l>4$. Then $3\leq m\leq n-3$. By Lemma 2.2,
 $d(T)$ $=d(T-v_{0})+d(T-v_{0}-v_{1})+d(T-v_{0}-v_{1}-v_{2}).$
 By Theorem 2.4,
-[…]
-  • $d(T-v_{0})=f(n-1)$.
+$\begin{aligned} d(T-v_{0}-v_{1}-v_{2})&\leq 2^{n-3-m}\cdot f(m)\leq 2^{n-6}\cdot f(3)=7\cdot 2^{n-6}\\ \end{aligned}.$
+We have found that there are two different cases of $d(T-v_{0})$. Let us discuss them separately below.
+Case 1: $d(T-v_{0})=f(n-1)$.
 By Theorem 2.4, $T-v_{0}\cong F_{n-1}$.
@@
 $=7\cdot 2^{n-4}+(n+2)\cdot 2^{\frac{n-4}{2}}$
-    $<h(n)=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}.$
+[…]
 This is contradictory.
@@
 This is contradictory.
-  • $d(T-v_{0})<f(n-1)$.
+Case 2: $d(T-v_{0})<f(n-1)$.
 Note that $T-v_{0}$ is connected and $d(T-v_{0})<f(n-1)$. By induction hypothesis, $d(T-v_{0})\leq h(n-1)$. Note that $T-v_{0}-v_{1}$ is connected. By Theorem 2.4, $d(T-v_{0}-v_{1})\leq f(n-2)$.
@@
 $=[2^{n-2}+(n+8)\cdot 2^{\frac{n-8}{2}}]+[2^{n-3}+(n+4)\cdot 2^{\frac{n-8}{2}}]+7\cdot 2^{n-6}$
-    $=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}-(2^{n-6}-n\cdot 2^{\frac{n-8}{2}})$
-    $<h(n)=2^{n-1}+(n+12)\cdot 2^{\frac{n-8}{2}}.$
-    The last inequality follows from the fact that $02^{n-6}-n\cdot 2^{\frac{n-8}{2}}>$ 0 when $n\geq 12$. This is contradictory.
+[…]
+[…]
+The last inequality follows from the fact that $2^{n-6}-n\cdot 2^{\frac{n-8}{2}}>0$ when $n\geq 12$. This is contradictory.
 Suppose that $n$ is odd ($n\geq 11$). By Lemma 2.1,
@@
 $\leq h(n-1)+f(n-2)+7\cdot 2^{n-6}$
-    $=[2^{n-2}+(n+11)\cdot 2^{\frac{n-9}{2}}]+[2^{n-3}+(n+1)\cdot 2^{\frac{n-7}{2}}]+7\cdot 2^{n-6}$
+[…]
 $=2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}-(2^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}})$
 $<h(n)=2^{n-1}+(n+9)\cdot 2^{\frac{n-7}{2}}.$
-    The last inequality follows from the fact that $02^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}}>$ 0 when $n\geq 11$. This is contradictory. Here, we have completed the proof of Claim $3$.
+The last inequality follows from the fact that $2^{n-6}-(n-5)\cdot 2^{\frac{n-9}{2}}>0$ when $n\geq 11$. This is contradictory. Here, we have completed the proof of Claim $3$.
 By Claim $3$, the component of $T-v_{2}$ that contains $v_{3}$ is a star $K_{1,x}$.
@@
 To complete the proof, we just need to show that $d(G)<h(n)$ if $c(G)=d\geq 2$.
-In fact, suppose $d\geq 2$. Since $G$ is connected, we can choose $d-1$ edges $u_{1}v_{1},u_{2}v_{2},\ldots,u_{d-1}v_{d-1}$ of $G$ such that the subgraph
+In fact, suppose $d\geq 2$. Since $G$ is connected, we can choose d1d-1 edges $u_{1}v_{1},u_{2}v_{2},\ldots,u_{d-1}v_{d-1}$ of $G$ such that the subgraph
 $G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-1}v_{d-1}:={G}^{\prime}$
@@
 $d(G)$ $\leq d(G-u_{1}v_{1})$
-[…]
+$\leq\cdots$
 $\leq d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})$
@@
 Hence, $d(G)=h(n)$ if and only if all of the above inequalities hold.
-By Theorem 2.5, $d({G}^{\prime})=h(n)$ is equivalent to ${G}^{\prime}\cong U_{n}$.
-By Lemma 2.3, $d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})=d({G}^{\prime})$ is equivalent to $u_{d-1},v_{d-1}$ are true twins in $G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2}$ It is follows that $u_{d-1},v_{d-1}$ are false twins in ${G}^{\prime}$.
+By Theorem 2.5, $d({G}^{\prime})=h(n)$ if and only if ${G}^{\prime}\cong U_{n}$.
+By Lemma 2.3, $d(G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2})=d({G}^{\prime})$ is equivalent to $u_{d-1}$ and $v_{d-1}$ are true twins in $G-u_{1}v_{1}-u_{2}v_{2}-\cdots-u_{d-2}v_{d-2}$.
+It follows that $u_{d-1}$ and $v_{d-1}$ are false twins in ${G}^{\prime}$.
 But this contradicts the fact that there are no false twins in ${G}^{\prime}\cong U_{n}$, where $U_{n}$ is defined in Theorem 2.5.
@@
 By Lemma 2.6, it is easy to see that
-$\begin{array}[]{llll}d(T_{2}^{9})<d(T_{1}^{9}),&d(T_{7}^{9})<d(T_{3}^{9}),&d(T_{8}^{9})<d(T_{5}^{9}),&d(T_{9}^{9})<d(T_{6}^{9}),\vspace{1.0ex}\\ d(T_{13}^{9})<d(T_{10}^{9}),&d(T_{19}^{9})<d(T_{1}^{9}),&d(T_{20}^{9})<d(T_{3}^{9}),&d(T_{21}^{9})<d(T_{5}^{9}),\vspace{1.0ex}\\ d(T_{23}^{9})<d(T_{4}^{9}),&d(T_{24}^{9})<d(T_{4}^{9}),&d(T_{25}^{9})<d(T_{6}^{9}),&d(T_{28}^{9})<d(T_{3}^{9}),\vspace{1.0ex}\\ d(T_{29}^{9})<d(T_{5}^{9}),&d(T_{30}^{9})<d(T_{22}^{9}),&d(T_{31}^{9})<d(T_{6}^{9}),&d(T_{32}^{9})<d(T_{10}^{9}),\vspace{1.0ex}\\ d(T_{33}^{9})<d(T_{22}^{9}),&d(T_{36}^{9})<d(T_{10}^{9}),&d(T_{38}^{9})<d(T_{27}^{9}),&d(T_{39}^{9})<d(T_{27}^{9}),\vspace{1.0ex}\\ d(T_{40}^{9})<d(T_{26}^{9}),&d(T_{41}^{9})<d(T_{27}^{9}),&d(T_{43}^{9})<d(T_{26}^{9}),&d(T_{44}^{9})<d(T_{4}^{9}),\vspace{1.0ex}\\ d(T_{45}^{9})<d(T_{6}^{9}),&d(T_{46}^{9})<d(T_{26}^{9}),&d(T_{47}^{9})<d(T_{6}^{9}).\end{array}$
+$\begin{array}[]{llll}d(T_{2}^{9})<d(T_{1}^{9}),&d(T_{7}^{9})<d(T_{3}^{9}),&d(T_{8}^{9})<d(T_{5}^{9}),&d(T_{9}^{9})<d(T_{6}^{9}),\vskip 4.30554pt\\ d(T_{13}^{9})<d(T_{10}^{9}),&d(T_{19}^{9})<d(T_{1}^{9}),&d(T_{20}^{9})<d(T_{3}^{9}),&d(T_{21}^{9})<d(T_{5}^{9}),\vskip 4.30554pt\\ d(T_{23}^{9})<d(T_{4}^{9}),&d(T_{24}^{9})<d(T_{4}^{9}),&d(T_{25}^{9})<d(T_{6}^{9}),&d(T_{28}^{9})<d(T_{3}^{9}),\vskip 4.30554pt\\ d(T_{29}^{9})<d(T_{5}^{9}),&d(T_{30}^{9})<d(T_{22}^{9}),&d(T_{31}^{9})<d(T_{6}^{9}),&d(T_{32}^{9})<d(T_{10}^{9}),\vskip 4.30554pt\\ d(T_{33}^{9})<d(T_{22}^{9}),&d(T_{36}^{9})<d(T_{10}^{9}),&d(T_{38}^{9})<d(T_{27}^{9}),&d(T_{39}^{9})<d(T_{27}^{9}),\vskip 4.30554pt\\ d(T_{40}^{9})<d(T_{26}^{9}),&d(T_{41}^{9})<d(T_{27}^{9}),&d(T_{43}^{9})<d(T_{26}^{9}),&d(T_{44}^{9})<d(T_{4}^{9}),\vskip 4.30554pt\\ d(T_{45}^{9})<d(T_{6}^{9}),&d(T_{46}^{9})<d(T_{26}^{9}),&d(T_{47}^{9})<d(T_{6}^{9}).\end{array}$
 By Corollary 2.7, $T_{12}^{9}$, $T_{15}^{9}$, $T_{16}^{9}$, $T_{17}^{9}$, $T_{18}^{9}$, $T_{34}^{9}$, $T_{35}^{9}$, $T_{37}^{9}$, $T_{42}^{9}$ are not the trees with the second-largest number of dissociation sets.
@@
 $=d(P_{2})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{1})\cdot d(P_{4})+d(P_{4})+d(P_{1})\cdot d(P_{3})$
-$=271<h(9);$
+[…]
 $d(T_{4}^{9})$ $=d(T_{4}^{9}-v_{0})+d(T_{4}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{4}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{3})\cdot d(P_{4})\cdot d(P_{1})+2\cdot d(P_{2})\cdot d(P_{3})+d(P_{1})\cdot d(P_{3})+d(P_{2})^{2}$
-$=268<h(9);$
+[…]
 $d(T_{5}^{9})$ $=d(T_{5}^{9}-v_{0})+d(T_{5}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{5}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{2})^{2}\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})+d(P_{1})\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})$
-$=280<h(9);$
+[…]
 $d(T_{6}^{9})$ $=d(T_{6}^{9}-v_{0})+d(T_{6}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{6}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{2})\cdot d(P_{3})^{2}+d(P_{1})\cdot d(P_{2})^{2}+d(P_{2})\cdot d(P_{1})^{2}\cdot 2+d(P_{2})^{2}$
-$=276<h(9);$
+[…]
 $d(T_{10}^{9})$ $=d(T_{10}^{9}-v_{0})+d(T_{10}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{10}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{2})^{2}\cdot d(P_{3})+d(P_{2})\cdot d(P_{1})^{2}\cdot 2+2\cdot d(P_{1})\cdot d(P_{2})+d(P_{1})^{3}$
-$=280<h(9);$
+[…]
 $d(T_{22}^{9})$ $=d(T_{22}^{9}-v_{0})+d(T_{22}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{22}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{4})^{2}+d(P_{1})^{2}\cdot d(P_{2})^{2}+2\cdot d(P_{1})^{2}\cdot d(P_{2})$
-$=265<h(9);$
+[…]
 $d(T_{26}^{9})$ $=d(T_{26}^{9}-v_{0})+d(T_{26}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{26}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{2})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{1})^{2}\cdot d(P_{3})+d(P_{1})\cdot d(P_{3})+d(P_{1})\cdot d(P_{2})$
-$=270<h(9);$
+[…]
 $d(T_{27}^{9})$ $=d(T_{27}^{9}-v_{0})+d(T_{27}^{9}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{27}^{9}-N[v_{0}]\cup N[u])$
 $=d(P_{2})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{1})\cdot d(P_{2})^{2}+d(P_{2})^{2}+d(P_{1})^{3}$
-$=280<h(9).$
+[…]
 Figure 8: Eight trees that may have the second-largest number of dissociation sets.
@@
 By Lemma 2.6, it is easy to see that
-$\begin{array}[]{llll}d(T_{2}^{10})<d(T_{1}^{10}),&d(T_{9}^{10})<d(T_{3}^{10}),&d(T_{10}^{10})<d(T_{5}^{10}),&d(T_{11}^{10})<d(T_{7}^{10}),\vspace{1.0ex}\\ d(T_{16}^{10})<d(T_{12}^{10}),&d(T_{17}^{10})<d(T_{13}^{10}),&d(T_{18}^{10})<d(T_{14}^{10}),&d(T_{27}^{10})<d(T_{1}^{10}),\vspace{1.0ex}\\ d(T_{28}^{10})<d(T_{3}^{10}),&d(T_{29}^{10})<d(T_{5}^{10}),&d(T_{31}^{10})<d(T_{4}^{10}),&d(T_{32}^{10})<d(T_{6}^{10}),\vspace{1.0ex}\\ d(T_{33}^{10})<d(T_{7}^{10}),&d(T_{36}^{10})<d(T_{4}^{10}),&d(T_{37}^{10})<d(T_{7}^{10}),&d(T_{39}^{10})<d(T_{8}^{10}),\vspace{1.0ex}\\ d(T_{44}^{10})<d(T_{3}^{10}),&d(T_{45}^{10})<d(T_{30}^{10}),&d(T_{46}^{10})<d(T_{5}^{10}),&d(T_{47}^{10})<d(T_{34}^{10}),\vspace{1.0ex}\\ d(T_{48}^{10})<d(T_{7}^{10}),&d(T_{49}^{10})<d(T_{35}^{10}),&d(T_{50}^{10})<d(T_{35}^{10}),&d(T_{51}^{10})<d(T_{12}^{10}),\vspace{1.0ex}\\ d(T_{52}^{10})<d(T_{38}^{10}),&d(T_{53}^{10})<d(T_{7}^{10}),&d(T_{54}^{10})<d(T_{41}^{10}),&d(T_{55}^{10})<d(T_{42}^{10}),\vspace{1.0ex}\\ d(T_{56}^{10})<d(T_{13}^{10}),&d(T_{57}^{10})<d(T_{43}^{10}),&d(T_{58}^{10})<d(T_{41}^{10}),&d(T_{60}^{10})<d(T_{14}^{10}),\vspace{1.0ex}\\ d(T_{63}^{10})<d(T_{12}^{10}),&d(T_{65}^{10})<d(T_{13}^{10}),&d(T_{67}^{10})<d(T_{59}^{10}),&d(T_{68}^{10})<d(T_{14}^{10}),\vspace{1.0ex}\\ d(T_{73}^{10})<d(T_{30}^{10}),&d(T_{74}^{10})<d(T_{35}^{10}),&d(T_{75}^{10})<d(T_{59}^{10}),&d(T_{76}^{10})<d(T_{43}^{10}),\vspace{1.0ex}\\ d(T_{77}^{10})<d(T_{41}^{10}),&d(T_{80}^{10})<d(T_{59}^{10}),&d(T_{83}^{10})<d(T_{8}^{10}),&d(T_{84}^{10})<d(T_{41}^{10}),\vspace{1.0ex}\\ d(T_{85}^{10})<d(T_{40}^{10}),&d(T_{86}^{10})<d(T_{42}^{10}),&d(T_{88}^{10})<d(T_{7}^{10}),&d(T_{89}^{10})<d(T_{34}^{10}),\vspace{1.0ex}\\ d(T_{90}^{10})<d(T_{38}^{10}),&d(T_{91}^{10})<d(T_{4}^{10}),&d(T_{92}^{10})<d(T_{6}^{10}),&d(T_{93}^{10})<d(T_{13}^{10}),\vspace{1.0ex}\\ d(T_{94}^{10})<d(T_{41}^{10}),&d(T_{95}^{10})<d(T_{7}^{10}),&d(T_{96}^{10})<d(T_{41}^{10}),&d(T_{97}^{10})<d(T_{42}^{10}),\vspace{1.0ex}\\ d(T_{98}^{10})<d(T_{87}^{10}),&d(T_{99}^{10})<d(T_{34}^{10}),&d(T_{100}^{10})<d(T_{38}^{10}),&d(T_{101}^{10})<d(T_{13}^{10}),\vspace{1.0ex}\\ d(T_{102}^{10})<d(T_{87}^{10}),&d(T_{103}^{10})<d(T_{41}^{10}),&d(T_{104}^{10})<d(T_{8}^{10}),&d(T_{105}^{10})<d(T_{40}^{10}).\end{array}$
+$\begin{array}[]{llll}d(T_{2}^{10})<d(T_{1}^{10}),&d(T_{9}^{10})<d(T_{3}^{10}),&d(T_{10}^{10})<d(T_{5}^{10}),&d(T_{11}^{10})<d(T_{7}^{10}),\vskip 4.30554pt\\ d(T_{16}^{10})<d(T_{12}^{10}),&d(T_{17}^{10})<d(T_{13}^{10}),&d(T_{18}^{10})<d(T_{14}^{10}),&d(T_{27}^{10})<d(T_{1}^{10}),\vskip 4.30554pt\\ d(T_{28}^{10})<d(T_{3}^{10}),&d(T_{29}^{10})<d(T_{5}^{10}),&d(T_{31}^{10})<d(T_{4}^{10}),&d(T_{32}^{10})<d(T_{6}^{10}),\vskip 4.30554pt\\ d(T_{33}^{10})<d(T_{7}^{10}),&d(T_{36}^{10})<d(T_{4}^{10}),&d(T_{37}^{10})<d(T_{7}^{10}),&d(T_{39}^{10})<d(T_{8}^{10}),\vskip 4.30554pt\\ d(T_{44}^{10})<d(T_{3}^{10}),&d(T_{45}^{10})<d(T_{30}^{10}),&d(T_{46}^{10})<d(T_{5}^{10}),&d(T_{47}^{10})<d(T_{34}^{10}),\vskip 4.30554pt\\ d(T_{48}^{10})<d(T_{7}^{10}),&d(T_{49}^{10})<d(T_{35}^{10}),&d(T_{50}^{10})<d(T_{35}^{10}),&d(T_{51}^{10})<d(T_{12}^{10}),\vskip 4.30554pt\\ d(T_{52}^{10})<d(T_{38}^{10}),&d(T_{53}^{10})<d(T_{7}^{10}),&d(T_{54}^{10})<d(T_{41}^{10}),&d(T_{55}^{10})<d(T_{42}^{10}),\vskip 4.30554pt\\ d(T_{56}^{10})<d(T_{13}^{10}),&d(T_{57}^{10})<d(T_{43}^{10}),&d(T_{58}^{10})<d(T_{41}^{10}),&d(T_{60}^{10})<d(T_{14}^{10}),\vskip 4.30554pt\\ d(T_{63}^{10})<d(T_{12}^{10}),&d(T_{65}^{10})<d(T_{13}^{10}),&d(T_{67}^{10})<d(T_{59}^{10}),&d(T_{68}^{10})<d(T_{14}^{10}),\vskip 4.30554pt\\ d(T_{73}^{10})<d(T_{30}^{10}),&d(T_{74}^{10})<d(T_{35}^{10}),&d(T_{75}^{10})<d(T_{59}^{10}),&d(T_{76}^{10})<d(T_{43}^{10}),\vskip 4.30554pt\\ d(T_{77}^{10})<d(T_{41}^{10}),&d(T_{80}^{10})<d(T_{59}^{10}),&d(T_{83}^{10})<d(T_{8}^{10}),&d(T_{84}^{10})<d(T_{41}^{10}),\vskip 4.30554pt\\ d(T_{85}^{10})<d(T_{40}^{10}),&d(T_{86}^{10})<d(T_{42}^{10}),&d(T_{88}^{10})<d(T_{7}^{10}),&d(T_{89}^{10})<d(T_{34}^{10}),\vskip 4.30554pt\\ d(T_{90}^{10})<d(T_{38}^{10}),&d(T_{91}^{10})<d(T_{4}^{10}),&d(T_{92}^{10})<d(T_{6}^{10}),&d(T_{93}^{10})<d(T_{13}^{10}),\vskip 4.30554pt\\ d(T_{94}^{10})<d(T_{41}^{10}),&d(T_{95}^{10})<d(T_{7}^{10}),&d(T_{96}^{10})<d(T_{41}^{10}),&d(T_{97}^{10})<d(T_{42}^{10}),\vskip 4.30554pt\\ d(T_{98}^{10})<d(T_{87}^{10}),&d(T_{99}^{10})<d(T_{34}^{10}),&d(T_{100}^{10})<d(T_{38}^{10}),&d(T_{101}^{10})<d(T_{13}^{10}),\vskip 4.30554pt\\ d(T_{102}^{10})<d(T_{87}^{10}),&d(T_{103}^{10})<d(T_{41}^{10}),&d(T_{104}^{10})<d(T_{8}^{10}),&d(T_{105}^{10})<d(T_{40}^{10}).\end{array}$
 Figure 9: When $n=10$, 106 structures of tree $T$.
@@
 $=d(P_{2})\cdot d(P_{1})\cdot d(P_{6})+2\cdot d(P_{1})\cdot d(P_{5})+d(P_{5})+d(P_{1})\cdot d(P_{4})$
-$=498<h(10);$
+[…]
 $d(T_{4}^{10})$ $=d(T_{4}^{10}-v_{0})+d(T_{4}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{4}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{3})\cdot d(P_{1})\cdot d(P_{5})+2\cdot d(P_{2})\cdot d(P_{4})+d(P_{1})\cdot d(P_{4})+d(P_{2})\cdot d(P_{3})$
-$=494<h(10);$
+[…]
 $d(T_{5}^{10})$ $=d(T_{5}^{10}-v_{0})+d(T_{5}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{5}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{2})^{2}\cdot d(P_{5})+d(P_{1})^{2}\cdot d(P_{4})+2\cdot d(P_{1})\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})$
-$=516<h(10);$
+[…]
 $d(T_{6}^{10})$ $=d(T_{6}^{10}-v_{0})+d(T_{6}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{6}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{4})^{2}\cdot d(P_{1})+d(P_{3})^{2}\cdot 2+2\cdot d(P_{2})\cdot d(P_{3})$
-$=492<h(10);$
+[…]
 $d(T_{7}^{10})$ $=d(T_{7}^{10}-v_{0})+d(T_{7}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{7}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{2})\cdot d(P_{7})+d(P_{1})\cdot d(P_{2})\cdot d(P_{4})+d(P_{2})\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})$
-$=508<h(10);$
+[…]
 $d(T_{8}^{10})$ $=d(T_{8}^{10}-v_{0})+d(T_{8}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{8}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{3})^{3}+d(P_{2})^{3}+3\cdot d(P_{1})\cdot d(P_{2})^{2}$
-$=503<h(10);$
+[…]
 $d(T_{12}^{10})$ $=d(T_{12}^{10}-v_{0})+d(T_{12}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{12}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{2})^{2}\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{3})\cdot 2+d(P_{1})\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})$
-$=516<h(10);$
+[…]
 $d(T_{13}^{10})$ $=d(T_{13}^{10}-v_{0})+d(T_{13}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{13}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{3})^{2}\cdot d(P_{2})+d(P_{2})^{2}\cdot d(P_{1})\cdot 2+d(P_{2})^{2}+d(P_{1})^{2}\cdot d(P_{2})\cdot 2$
-$=504<h(10);$
+[…]
 $d(T_{14}^{10})$ $=d(T_{14}^{10}-v_{0})+d(T_{14}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{14}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{2})^{3}\cdot d(P_{3})+d(P_{1})^{3}\cdot d(P_{2})+d(P_{1})^{2}\cdot d(P_{2})\cdot 3+d(P_{1})^{4}$
-$=544<h(10);$
+[…]
 Figure 10: Nineteen trees that may have the second-largest number of dissociation sets.
 $d(T_{30}^{10})$ $=d(T_{30}^{10}-v_{0})+d(T_{30}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{30}^{10}-N[v_{0}]\cup N[u])$
-$=d(P_{4})\cdot 23+d(P_{1})\cdot d(P_{2})\cdot d(P_{4})+d(P_{1})\cdot d(P_{4})+d(P_{1})^{2}\cdot d(P_{2})^{2}$
-$=493<h(10);$
+[…]
+[…]
 $d(T_{34}^{10})$ $=d(T_{34}^{10}-v_{0})+d(T_{34}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{34}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{4})\cdot d(P_{5})+d(P_{1})^{2}\cdot d(P_{2})\cdot d(P_{3})+d(P_{1})^{2}\cdot d(P_{3})+d(P_{1})\cdot d(P_{2})^{2}$
-$=484<h(10);$
+[…]
 $d(T_{35}^{10})$ $=d(T_{35}^{10}-v_{0})+d(T_{35}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{35}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{4})\cdot d(P_{5})+d(P_{1})\cdot d(P_{2})^{3}+d(P_{1})\cdot d(P_{2})^{2}+d(P_{1})^{3}\cdot d(P_{2})$
-$=504<h(10);$
+[…]
 $d(T_{38}^{10})$ $=d(T_{38}^{10}-v_{0})+d(T_{38}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{38}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{4})^{2}+d(P_{1})\cdot d(P_{2})\cdot d(P_{3})\cdot 2+d(P_{1})\cdot d(P_{3})+d(P_{2})^{2}\cdot d(P_{1})$
-$=496<h(10);$
+[…]
 $d(T_{40}^{10})$ $=d(T_{40}^{10}-v_{0})+d(T_{40}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{40}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{3})\cdot d(P_{5})+d(P_{1})\cdot d(P_{2})\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{3})+d(P_{2})^{2}$
-$=492<h(10);$
+[…]
 $d(T_{41}^{10})$ $=d(T_{41}^{10}-v_{0})+d(T_{41}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{41}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{2})\cdot d(P_{6})+d(P_{1})\cdot d(P_{2})\cdot d(P_{3})\cdot 2+d(P_{2})\cdot d(P_{3})+d(P_{1})^{2}\cdot d(P_{2})$
-$=508<h(10);$
+[…]
 $d(T_{42}^{10})$ $=d(T_{42}^{10}-v_{0})+d(T_{42}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{42}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{2})^{2}\cdot d(P_{5})+d(P_{1})^{3}\cdot d(P_{3})+d(P_{1})^{2}\cdot d(P_{3})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})$
-$=512<h(10);$
+[…]
 $d(T_{43}^{10})$ $=d(T_{43}^{10}-v_{0})+d(T_{43}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{43}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{2})^{2}\cdot d(P_{5})+d(P_{1})^{2}\cdot d(P_{2})^{2}+d(P_{1})\cdot d(P_{2})^{2}\cdot 2+d(P_{1})^{4}$
-$=528<h(10);$
+[…]
 $d(T_{59}^{10})$ $=d(T_{59}^{10}-v_{0})+d(T_{59}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{59}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{2})^{2}\cdot d(P_{4})+d(P_{1})^{3}\cdot d(P_{2})\cdot 2+d(P_{1})^{2}\cdot d(P_{2})\cdot 2+d(P_{1})^{3}$
-$=520<h(10);$
+[…]
 $d(T_{87}^{10})$ $=d(T_{87}^{10}-v_{0})+d(T_{87}^{10}-N[v_{0}])+\sum_{u\in N(v_{0})}d(T_{87}^{10}-N[v_{0}]\cup N[u])$
 $=d(P_{1})\cdot d(P_{4})^{2}+d(P_{1})^{2}\cdot d(P_{2})^{2}\cdot 2+d(P_{1})^{2}\cdot d(P_{2})\cdot 2$
-$=498<h(10).$
 […]
+[…]